21x^2+2x-3= (7x+3)(3x-1)
Check: (7x+3)(3x-1)=7x(3x)+7x(-1)+3(3x)+3(-1)=21x^2-7x+9x-3-21^2+2x-3

Answer by mattiewal(1)

21x^2+2x-3= (7x+3)(3x-1)
Check: (7x+3)(3x-1)=7x(3x)+7x(-1)+3(3x)+3(-1)=21x^2-7x+9x-3-21^2+2x-3